OpenStudy (ilovecomputers1):

What is the equation in point-slope form of the line passing through (0, 5) and (−2, 11)? y − 5 = −3(x + 2) y − 5 = 3(x + 2) y − 11 = −3(x − 2) y − 11 = −3(x + 2) I got y − 11 = −3(x + 2)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 0}}\quad ,&{\color{blue}{ 5}})\quad &({\color{red}{ -2}}\quad ,&{\color{blue}{ 11}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values }\\ \qquad \uparrow\\ \textit{point-slope form} \)

OpenStudy (ilovecomputers1):

So then I get y-5=-3(x-0)? @jdoe0001

OpenStudy (jdoe0001):

yeap

OpenStudy (ilovecomputers1):

Alright thank you. That's not one of my answer choices though, so one of them must be a mistake.. I think I know which one to put down though. Thank you very much.

OpenStudy (jdoe0001):

right, because likely you'd need to distribute the -3

OpenStudy (jdoe0001):

so y-5 = -3x

OpenStudy (jdoe0001):

hmm I see your choices... hmm